OpenStudy (anonymous):

WILL FAN AND MEDAL 1/5x+3(x-2)=26

OpenStudy (anonymous):

@YoungStudier

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (anonymous):

i started with the parenthesis 1/5x+3x-6=26

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@FortyTheRapper

OpenStudy (anonymous):

@IsabellaVH

OpenStudy (skullpatrol):

Now add 6 to both sides please.

OpenStudy (anonymous):

1/5x+3x= 32

OpenStudy (anonymous):

combining like terms is a challenge .....i hate fractions

OpenStudy (skullpatrol):

Hate is such a strong word. Try your best :-) (1/5 + 3)x

OpenStudy (anonymous):

Fine i strongly dislike fractions 3 1/5x =32

OpenStudy (nerdsarecool):

What do we need to find out

OpenStudy (skullpatrol):

3 1/5 = what as an improper fraction?

OpenStudy (anonymous):

\[3 \frac{ 1 }{ 5 }\] no as a mixed fraction

OpenStudy (crystal_bliss):

Simplify 1/5x to x/5 x/5+3(x−2)=26 2 Multiply both sides by 5 x+15(x−2)=130 3 Expand x+15x−30=130 4 Simplify x+15x−30 to 16x−30 16x−30=130 5 Add 30 to both sides 16x=130+30 6 Simplify 130+30 to 160 16x=160 7 Divide both sides by 16 x=160/16 8 Simplify 160/16 to 10 x=10

OpenStudy (anonymous):